# Fine-Grained Vision-Language Modeling for Multimodal Training Assistants in Augmented Reality

### v1

@@ Abstract
VLM are essential for enabling AI-powered smart assistants to interpret and reason in multimodal environments. However, their application in augmented reality (AR) training remains largely unexplored. In this work, we introduce a comprehensive dataset tailored for AR training, featuring systematized vision-language tasks, and evaluate nine state-of-the-art VLMs on it. Our results reveal that even advanced models, including GPT-4o, struggle with fine-grained assembly tasks, achieving a maximum F1 score of just 40.54% on state detection. These findings highlight the demand for enhanced datasets, benchmarks, and further research to improve fine-grained vision-language alignment. Beyond technical contributions, our work has broader social implications, particularly in empowering blind and visually impaired users with equitable access to AI-driven learning opportunities. We provide all related resources, including the dataset, source code, and evaluation results, to support the research community.
1 Introduction
Training assistants are in high demand in industrial assembly tasks, including furniture production (You et al., 2022), car manufacturing (Bellalouna et al., 2020), and industrial product assembly (Funk et al., 2017). As these tasks often require high levels of precision and cognitive effort, augmented reality (AR) has emerged as a promising technology to support workers, providing a multimodal, interactive, user-centric environment (Hou & Wang, 2013; Botto et al., 2020; Dalim et al., 2020). AR training systems have been extensively studied for their ability to enhance user experience and task efficiency (Büttner et al., 2020; Blattgerste et al., 2021). However, most existing research focuses on interaction design and usability studies, rather than the development of intelligent, instruction-following assistants that can understand and process complex manual-based assembly tasks in AR environments.
@@ 1 Introduction
VLM offer a powerful solution for bridging visual and textual modalities by effectively interpreting and integrating visual inputs with corresponding textual instructions. Traditional computer vision studies have explored various purely visual tasks, such as action segmentation, recognition, anticipation, and object and state detection (Damen et al., 2018; Miech et al., 2020; Wang et al., 2023b; Gao et al., 2024). Some studies focus on analyzing procedural activities in videos without fixed instructions (Tang et al., 2019; Sener et al., 2022), while others incorporate textual instructions to enhance comprehension. These approaches include using instructional queries for response generation(Li et al., 2023a; Padmakumar et al., 2021), narration-video alignment(Zhukov et al., 2019), and text-to-video retrieval(Miech et al., 2019). However, none of these studies explicitly address instruction-following in manual assembly tasks, where models must accurately interpret step-by-step procedural instructions and track fine-grained object states within a structured pipeline. The recent advancements in large language models and vision-language models have made multimodal reasoning more accessible, with models such as GPT-4(Achiam et al., 2023), LLaVa (Liu et al., 2023), QWEN-VL (Bai et al., 2023), BLIP2 (Li et al., 2023c), and MiniGPT-v2 (Chen et al., 2023). The core idea is to excel in multimodal tasks by integrating visual encoders with fundamental LLMs. Some early-stage explorations have attempted to apply VLMs in AR to align visual contexts with language representations to showcase the capabilities of LLM agents (Pei et al., 2024). However, these efforts have primarily focused on coarse-grained visual features, such as detecting objects and their coordinates, rather than understanding complex procedural assembly tasks in detail. A key challenge in AR-based assembly training is the fine-grained interpretation of object states and spatial relationships. Figure 1 illustrates this limitation: when prompted with an image and text instructions for LEGO assembly, GPT-4o fails to correctly identify an incorrect assembly state, where a newly added LEGO part (in green) is misplaced. This example highlights a crucial weakness – current VLMs struggle with precise state detection and fine-grained assembly verification, making them unreliable for industrial AR training scenarios.
Figure 1: An example of incorrect state detection by GPT-4o during LEGO assembly. Given an image and text instructions, the model fails to accurately recognize an incorrectly placed LEGO part (highlighted in green), demonstrating the challenge of fine-grained vision-language alignment in assembly tasks.
To address these challenges, we introduce a comprehensive vision-language benchmark and dataset designed for AR-based assembly training. Specifically, we chose LEGO brick assembly in AR as a simplified simulation and created a synthetic multimodal dataset to address vision-language challenges, i.e., complex scene understanding, object detection, and state detection. We developed a unified, fine-grained VLM architecture capable of handling a broad range of vision-language tasks in AR and validated the feasibility of nine prevailing models as benchmarks. Notably, this research is designed for various manual-guided assembly assistance in a multimodal environment, though we specifically demonstrated it using LEGO assembly. We list contributions below: (1) Developing a unified VLMs-based architecture to benchmark fine-grained instruction-following capabilities in AR training. (2) Evaluating nine prevailing VLMs, both without and with fine-tuning, on a synthetic multimodal dataset tailored for AR-based LEGO assembly. (3) Creating and releasing a dataset containing systematic vision-language tasks for AR training, along with a method for synthesizing high-quality multimodal data.
2 Fine-grained Vision-Language Modeling
2.1 Task definition
As illustrated in Figure 2, we investigate manual-guided LEGO assembly in AR and define the following three vision-language tasks: (T1) Scene Understanding. Given the current step’s image $V_{t}$ and the task-specific query $Q_{t}$, the model outputs a response containing the scene description for the current assembly step $I_{t}$. $Q_{t}$ is the task description contextualized with the previous step’s textual instruction $I_{t-1}$. It assesses the model’s ability to accurately describe the current assembly step from an instruction manual, including but not limited to generating object representations, their properties, and the assembly procedure. (T2) Object Detection. Given a seeking step’s image $V_{t^{\prime}}$ ($t^{\prime}=t-1$) and task-specific query $Q_{t^{\prime}}$, the model outputs a response containing the positional text $P_{t^{\prime}}$, which includes identified object and its coordinate formatted as “<Object> <Xleft> <Ytop> <Xright> <Ybottom>”. The query $Q_{t^{\prime}}$ is the task description contextualized with the corresponding textual instruction $I_{t^{\prime}}$. It assesses the model’s ability to accurately identify and locate objects in the scene, ensuring they are positioned correctly for the next assembly action. (T3) State Detection. Given an assembly step’s image, either the original manual image $V_{t}$ or a generated negative sample $\tilde{V}_{t}$, and task-specific query $Q_{t}$, the model outputs a response indicating a correct or incorrect state. It evaluates the model’s ability to verify the accuracy of the assembly progression, determining if the assembly action has been correctly completed. These tasks have been crafted to rigorously test and demonstrate the capabilities of VLMs in an AR training, focusing on their ability to understand and interpret complex multimodal inputs.
2.2 One-for-all architecture
Figure 3: The architecture for one-for-all vision-language modeling. It integrates a vision encoder, a large language model, and a vision-language projector as core functional modules, along with a task-specific query for task adaptation.
3 Dataset Creation
3.1 Manual Crawling and Scene Data Matching
This subsection outlines the procedure for creating the dataset for task T1. First, we collected 65 instruction manuals from the official LEGO website [^1] https://legoaudioinstructions.com/instructions, designed to help blind and visually impaired users assemble LEGO sets accurately. Each instruction manual comprises several elements: (1) step-by-step textual instructions; (2) corresponding image for each step; and (3) tags provided by LEGO include instruction and image tags. Then, we split a full manual into several assembly sessions, each session includes two types of steps: object seeking and object assembly. Each seeking step is followed by an assembly step , forming a pair. Next, we iterated through all sessions to generate the scene understanding dataset, $D_{T1}=\{(Q_{t},V_{t},A_{t})\}|_{t=0}^{|D_{T1}|}$, where each element is a triplet of (query, image, text), and the query $Q_{t}$ is constructed by filling the query template with the task token “[grounding]” and the previous instruction $I_{t-1}$ as context.
3.2 Object Position Inference
This subsection outlines the procedure for creating the dataset for task T2, as illustrated in Figure 6.
First, we iterated through all object-seeking steps, where users are asked to find specific objects. Second, we iterated all images to generate the object detection dataset, $D_{T2}=\{(Q_{t^{\prime}},\tilde{V}_{t^{\prime}},A_{t^{\prime}})\}|_{t^{\prime}=0}^{|D_{T2}|}$, where each element is a triplet of (query, image, text) for a seeking step $t^{\prime}=t-1$. The query $Q_{t^{\prime}}$ is constructed by filling the query template with the task token “[object]” and the current instruction $I_{t^{\prime}}$. The response $A_{t^{\prime}}$ contains a positional text $\tilde{P}_{t^{\prime}}$ formatted as “<Object> <Xleft> <Ytop> <Xright> <Ybottom>”. This is initially generated by querying the image using MiniGPT-v2 (Chen et al., 2023), the state-of-the-art model at the time of this work. The composite image $\tilde{V}_{t^{\prime}}$ is created by combining the current image ${V}_{t^{\prime}}$ with three randomly sampled images. Last, the initial coordinates are adjusted to fit the composite image.
3.3 Variant State Generation
This subsection outlines the procedure for creating the dataset for task T3, as illustrated in Figure 7. First, we iterated through all object-assembly steps, along with their corresponding previous steps, where users are trained to assemble the parts identified in the prior steps. Second, we conducted part segmentation by detecting the boundary color and segmenting the part to be assembled. In each assembly session, the objects to be assembled in the current step are tagged in the previous step, with their boundaries highlighted in distinct colors that vary between sessions. The highlighted colors are selected based on predominant color detection using K-means clustering [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html, followed by Hue filtering (Chu et al., 2014) and manual correction for accuracy. Third, we added bounding boxes to the segmented objects. Fourth, we applied a natural perturbation to the part to be assembled by randomly shifting it within the background box, ensuring it moves by at least 5%. We treated each state from the manual as a positive sample and generated three variant states as negative samples. Then, we constructed the state detection dataset, $D_{T3}=\{(Q_{t},V_{t},A_{t},\tilde{V}_{t},\tilde{A}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}},\tilde{A}_{t}^{{}^{\prime\prime}})\}|_{t=0}^{|D_{T3}|}$, where the query $Q_{t}$ is constructed by filling the query template with the task token “[state]”, the current instruction $I_{t}$, and the previous instruction $I_{t-1}$ as context. The variant states $\{\tilde{V}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}}\}$ are natural pertubations of $V_{t}$. The responses $A_{t}$ and $\{\tilde{A}_{t},\tilde{A}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime\prime}}\}$ indicate the correct and incorrect statuses of the assembly state, respectively.
3.4 Dataset Statistics
Manual Session Scene Object State Step Sample
Overall Identification Assembly Overall T1 T2 T3
65 397 5,612 4,784 2,716 10,428 4,814 5,614 35,612 5,612 19,136 10,864
Table 1: Statistics of LEGO-VLM dataset.
4 Experimental Setup
4.1 Benchmarks
We benchmark the nine prevailing VLMs: (1) mPLUG-OWL2 (Ye et al., 2024)replaces attention with a modality adapter in a LLM decoder. (2) BLIP2 (Li et al., 2023c)uses a two-stage pre-trained Q-Former between an image encoder and a LLM to bridge the vision-language modality gap. (3) LLaVa (Liu et al., 2023)combines a visual CLIP encoder and a language decoder Vicuna for general-purpose visual language understanding. (4) Qwen-VL (Bai et al., 2023)connects a LLM with a visual encoder using position-aware vision-language adapter towards fine-grained visual understanding. (5) InstructBLIP (Dai et al., 2023)explores general-purpose vision-language instruction tuning based on the pretrained BLIP2. (6) MiniGPT-v2 (Chen et al., 2023)links a frozen ViT visual encoder with Llama-2 via a projection layer, applicable for diverse tasks via task-specific multimodal instructions. (7) Otter (Li et al., 2023a)is tuned on the OpenFlamingo, conditioning the language model on images for multi-modal perception and reasoning. (8) MiniGPT-4 (Zhu et al., 2023)integrates a frozen visual ViT&Q-Former encoder and Vicuna via a projection layer, unlocking advanced multimodal capabilities like GPT-4. (9) GPT-4o [^3] https://openai.com/index/hello-gpt-4o/is a widely used commercial model accessible via APIs.
4.2 Evaluation Metrics
[…]
5 Outcomes
5.1 Benchmark Results
@@
mPLUG-OWL2 Ye et al. (2024) 23.16 29.70 25.23 32.75 3.04 8.46 77.94 93.05 14.23 34.88 36.82 15.00 63.41 35.00
BLIP2 Li et al. (2023c) 27.85 32.65 29.55 40.35 6.40 12.50 77.35 84.48 34.25 40.57 24.71 28.50 100.00 94.87
LlaVA Liu et al. (2023) 32.10 35.88 13.63 42.97 0.99 17.73 54.39 72.32 $\emptyset$ 60.98- 25.04 30.00 99.41 50.00
QWEN-VL Bai et al. (2023) 34.84 37.00 29.57 39.08 5.19 13.13 78.56 89.15 25.60 30.08 39.77 39.53 97.99 96.82
InstructBLIP Dai et al. (2023) 32.85 32.85- 29.20 36.87 4.91 11.52 79.76 98.16 $\emptyset$ 47.20- 0.00 0.02 1.22 0.00
MiniGPT-v2 Chen et al. (2023) 33.06 37.52 34.72 32.56 9.81 8.28 84.95 85.91 26.98 25.94 36.76 38.64 60.42 80.72
Otter Li et al. (2023a) 11.49 12.12 1.28 72.39 $\emptyset$ 35.33 75.88
MiniGPT-4 Zhu et al. (2023) 34.11 15.05 1.88 87.09 30.20 37.19 67.37
GPT-4o 44footnotemark: 4 25.81 18.67 2.00 66.67 21.68 40.54 43.06
Table 2: Benchmarking vision-language models on LEGO-VLM dataset, without (/wo) and with (/w) PEFT using LoRA. The bold font indicates the highest score in each column. Symbols $\uparrow$ and $\downarrow$ denote that higher and lower values are better, respectively. Symbol “–” indicates the model is not applicable for fine-tuning. Symbol $\emptyset$” denotes a meaningless zero as the model fails to generate output as instructed. The superscripts “”, “”, and “-” indicate an increase, decrease, or inapplicability in the evaluation score after fine-tuning, respectively.
@@
5.2 Data Quality Assessment
@@
To ensure the quality of the generated data for tasks T2 and T3, we sampled 100 data points and conducted quality assessments for each task. We added bounding boxes based on the coordinates and asked three annotators to evaluate the generated data quality, focusing on coordinates, entities, and negative samples, based on the following criteria: (1) Entity disambiguitymeasures how clearly a target entity is relevant to the scene in T2, with scores of 0, 1, and 2 indicating low, medium, and high disambiguation, respectively. (2) Boundary precisionmeasures how accurately a bounding box encloses a target object in T2, with scores of 0, 1, and 2 indicating low, medium, and high precision, respectively. (3) State relevancemeasures whether the generated parts in an image are a relevant variation of the original image in T3. (4) State identifiabilitymeasures whether the generated parts in an image are a recognizable variation of the original image in T3.
Figure 4: Results of the quality assessment.
As shown in Figure 4, the average entity disambiguation and boundary precision scores are 1.74 and 1.25 out of 2 with the Kappa values [^5] https://www.statsmodels.org/dev/generated/statsmodels.stats.inter_rater.fleiss_kappa.html of 0.60 and 0.61, respectively. This suggests that annotators reached a moderate agreement on the clarity of the evaluated objects in relation to the corresponding scenes. The average state relevance and identifiability scores are 0.93 and 0.74 out of 1.00, respectively, with the Kappa scores 0.81 and 0.59. This indicates that annotators reached near-perfect agreement on the relevance of the generated states to the original while showing moderate agreement on the recognizable variation of those states. Besides, the first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria. This ensures the overall quality of the generated states.
5.3 Case Study
Text Image
@@ 5.3 Case Study
Query [grounding] After I find 1 transparent mailbox front 2x2. Tell me what to do for the current situation. [Uncaptioned image]
@@
Reference Insert the two tabs at the bottom of the mailbox casing into the two small notches on the mailbox front.
mPLUG-OWL2 Place the mailbox front 2x2 next to the block on the table.
Fine-tuned Put it on the rightmost edge of the high part of the structure.
@@
BLIP2 Put the mailbox front 2x2 horizontally on the table, clasp to the back. Put the mailbox back 2x2 vertically on top, slope to the left.
Fine-tuned Place the mailbox front 2x2 on the left side of the structure to complete the assembly.
@@
LlaVA In the image, there is a small Lego-like object with a red liquid inside of it. The object is placed in a blue sky, and it appears to be a 3D model. To create a transparent mailbox, you need to find a transparent mailbox front that is 2x2 in size. Once you find it, you can use it as the base for your transparent mailbox. You can then place the small Lego-like object inside the mailbox to complete the scene.
Fine-tuned Put it on the mailbox.
@@
QWEN-VL Put the transparent mailbox front 2x2 on the right side of the white arch.
Fine-tuned Put it on the table.
InstructBlip Place it horizontally on the table, clasp to the back.
Fine-tuned Put the mailbox front on the left side of the mailbox base.
@@
MiniGPT-v2 Sure! Here’s what you need to do with the pieces you have collected: 1. Place the transparent mailbox front 2x2 on top of the plate 1x2, aligned with the short side of the plate. Now you have completed the current situation. Let me know if you need help with anything else!
Fine-tuned Put the mailbox front 2x2 horizontally in front of the previous piece.
@@
Otter In view of the reasoning ‘A mailbox is small and square with a see through window.’, the optimal answer should be: build structure.
@@
MiniGPT-4 Attach the transparent mailbox front to the base structure, aligning the tabs with the notches.

### v2

# LEGO Co-builder: Exploring Fine-Grained Vision-Language Modeling for Multimodal LEGO Assembly Assistants

@@ Abstract
Vision-language models (VLMs) are facing the challenges of understanding and following multimodal assembly instructions, particularly when fine-grained spatial reasoning and precise object state detection are required. In this work, we explore LEGO Co-builder, a hybrid benchmark combining real-world LEGO assembly logic with programmatically generated multimodal scenes. The dataset captures stepwise visual states and procedural instructions, allowing controlled evaluation of instruction-following, object detection, and state detection. We introduce a unified framework and assess leading VLMs such as GPT-4o, Gemini, and Qwen-VL, under zero-shot and fine-tuned settings. Our results reveal that even advanced models like GPT-4o struggle with fine-grained assembly tasks, with a maximum F1 score of just 40.54% on state detection, highlighting gaps in fine-grained visual understanding. We release the benchmark, codebase, and generation pipeline to support future research on multimodal assembly assistants grounded in real-world workflows.
Introduction
Multimodal instruction-following assistants are gaining increasing relevance in domains requiring precise procedural understanding, such as furniture construction (You et al. 2022), automotive manufacturing (Bellalouna et al. 2020), and industrial product assembly (Funk et al. 2017). These tasks demand step-by-step reasoning, spatial awareness, and accurate interpretation of visual and textual instructions—capabilities that current AI systems still struggle to reliably deliver.
Figure 1: An example of incorrect state detection by GPT-4o during LEGO assembly. Given an image and text instructions, the model fails to accurately recognize an incorrectly placed LEGO part (highlighted in green), demonstrating the challenge of fine-grained vision-language alignment.
Traditional computer vision research has tackled various purely visual tasks, including action segmentation, recognition, anticipation, and object state detection (Damen et al. 2018; Miech et al. 2020; Wang et al. 2023b; Gao et al. 2024). Procedural activity understanding has also been explored using video datasets with or without textual narration (Tang et al. 2019; Zhukov et al. 2019; Li et al. 2023a; Sener et al. 2022). Some methods incorporate textual alignment, such as narration-video grounding or instruction-conditioned generation (Padmakumar et al. 2021; Miech et al. 2019), but often rely on separate architectures for each task type.
@@ Introduction
VLM offer a promising solution by bridging visual and textual modalities, enabling models to align procedural language with visual observations. Recent advances in large-scale models such as GPT-4o (Achiam et al. 2023), LLaVa (Liu et al. 2023), Qwen-VL (Bai et al. 2023), BLIP-2 (Li et al. 2023c), MiniGPT-v2 (Chen et al. 2023), and more recent reasoning-augmented variants like VLM-R1 (Shen et al. 2025), GLM-4.1v-thinking (Hong et al. 2025), and Gemini2.5 (Comanici et al. 2025), have demonstrated strong performance on a range of general-purpose multimodal tasks. However, most existing benchmarks focus on coarse-grained capabilities such as object recognition, captioning, and visual question answering, and do not evaluate the detailed, procedural understanding required for step-by-step manual assembly. While recent VLMs research (Laurençon et al. 2024; Cheng et al. 2024; Chen et al. 2024a) aim to unify vision-language reasoning under a single architecture, challenges remain—particularly for tasks that require dense and fine-grained perception (Wei et al. 2024; Rahmanzadehgervi et al. 2024; Chen et al. 2024b).
In this work, we address the need for instruction-following benchmarks that evaluate a model’s ability to reason over both visual and linguistic input in detail. We focus on LEGO brick assembly as a representative manual task that is richly structured, visually complex, and procedurally grounded. Figure 1 illustrates the limitations of current models in such tasks. When prompted with a visual scene and corresponding instruction, GPT-4o fails to detect a misplaced LEGO piece, highlighting the challenge of fine-grained visual scene understanding and object state recognition. This motivates the need for benchmarks that go beyond high-level visual understanding and evaluate precise spatial compliance with procedural steps.
To address this gap, we present LEGO Co-builder, a benchmark and dataset designed to evaluate fine-grained instruction-following capabilities in LEGO assembly tasks. Our dataset includes multimodal sequences consisting of visual snapshots, object states, and procedural steps derived from human-crafted manuals. We introduce a unified task formulation and evaluate nine leading VLMs in both zero-shot and fine-tuned settings. Our contributions are summarized as follows: (1) We develop a unified vision-language architecture to benchmark fine-grained, multimodal, instruction-following capabilities in procedural manual-guided assembly tasks. (2) We evaluate prevailing VLMs on a dataset of LEGO assembly sequences with grounded visual and textual supervision, under both zero-shot and fine-tuned settings. (3) We release the dataset, benchmarking code, and a modular synthetic data generation pipeline to support future research in multimodal instruction-following.
Societal Impact. This work advances the development of multimodal AI assistants that can transform how people learn and perform complex physical tasks. By enabling fine-grained vision-language understanding, our benchmark supports the creation of multimodal educational tools that combine visual input, language, and potentially augmented reality (AR) to enhance hands-on learning experiences. These systems can improve training in domains like education and industrial training. Moreover, the ability to interpret and verify procedural steps visually opens up promising avenues for assistive technologies—particularly for blind or visually impaired learners—by providing real-time, AI-driven guidance through tasks that were previously inaccessible. In this way, our research supports both innovation in teaching modalities and greater inclusivity in skill development.
Fine-Grained Vision-Language Modeling
Task Definition
We investigate manual-guided LEGO assembly and define the following three vision-language tasks (See in Figure 2):
  1. (T1)
    Scene Understanding. Given the current step’s image $V_{t}$ and the task-specific query $Q_{t}$, the model outputs a response containing the scene description for the current assembly step $I_{t}$. $Q_{t}$ is the task description contextualized with the previous step’s textual instruction $I_{t-1}$. It assesses the model’s ability to accurately describe the current assembly step from an instruction manual, including but not limited to generating object representations, their properties, and the assembly procedure.
  2. (T2)
    Object Detection. Given a seeking step’s image $V_{t^{\prime}}$ ($t^{\prime}=t-1$) and task-specific query $Q_{t^{\prime}}$, the model outputs a response containing the positional text $P_{t^{\prime}}$, which includes identified object and its coordinate formatted as “[Object] [Xleft] [Ytop] [Xright] [Ybottom]”. The query $Q_{t^{\prime}}$ is the task description contextualized with the corresponding textual instruction $I_{t^{\prime}}$. It assesses the model’s ability to accurately identify and locate objects in the scene, ensuring they are positioned correctly for the next assembly action.
  3. (T3)
    State Detection. Given an assembly step’s image, either the original manual image $V_{t}$ or a generated negative sample $\tilde{V}_{t}$, and task-specific query $Q_{t}$, the model outputs a response indicating a correct or incorrect state. It evaluates the model’s ability to verify the accuracy of the assembly progression, determining if the assembly action has been correctly completed.
These tasks have been crafted to rigorously test and demonstrate the capabilities of VLMs in LEGO Assembly, focusing on their ability to understand and interpret complex multimodal inputs.
One-for-all Vision-Language Architecture
Figure 3: The architecture for one-for-all vision-language modeling. It integrates a vision encoder, a large language model, and a vision-language projector as core functional modules, with a task-specific query for task adaptation.
Dataset Creation
Manual Crawling and Scene Data Matching
We outline the procedure for creating the dataset for task T1. First, we collected 65 official LEGO instruction manuals (LEGO 2024), designed to help blind and visually impaired users assemble LEGO sets accurately. Each instruction manual comprises several elements: (1) step-by-step textual instructions; (2) corresponding image for each step; and (3) tags provided by LEGO include instruction and image tags. Then, we split a full manual into several assembly sessions, each session includes two types of steps: object seeking and object assembly. Each seeking step is followed by an assembly step , forming a pair. Next, we iterated through all sessions to generate the scene understanding dataset, $D_{T1}=\{(Q_{t},V_{t},A_{t})\}|_{t=0}^{|D_{T1}|}$, where each element is a triplet of (query, image, text), and the query $Q_{t}$ is constructed by filling the query template with the task token “[grounding]” and the previous instruction $I_{t-1}$ as context.
Object Position Inference
This subsection outlines the procedure for creating the dataset for task T2. First, we iterated through all object-seeking steps, where users are asked to find specific objects. Second, we iterated all images to generate the object detection dataset, $D_{T2}=\{(Q_{t^{\prime}},\tilde{V}_{t^{\prime}},A_{t^{\prime}})\}|_{t^{\prime}=0}^{|D_{T2}|}$, where each element is a triplet of (query, image, text) for a seeking step $t^{\prime}=t-1$. The query $Q_{t^{\prime}}$ is constructed by filling the query template with the task token “[object]” and the current instruction $I_{t^{\prime}}$. The response $A_{t^{\prime}}$ contains a positional text $\tilde{P}_{t^{\prime}}$ formatted as “[Object] [Xleft] [Ytop] [Xright] [Ybottom]”. This is initially generated by querying the image using MiniGPT-v2 (Chen et al. 2023), the state-of-the-art model at the time of this work. The composite image $\tilde{V}_{t^{\prime}}$ is created by combining the current image ${V}_{t^{\prime}}$ with three randomly sampled images. Last, the initial coordinates are adjusted to fit the composite image.
Variant State Generation
This subsection outlines the procedure for creating the dataset for task T3. First, we iterated through all object-assembly steps, along with their corresponding previous steps, where users are trained to assemble the parts identified in the prior steps. Second, we conducted part segmentation by detecting the boundary color and segmenting the part to be assembled. In each assembly session, the objects to be assembled in the current step are tagged in the previous step, with their boundaries highlighted in distinct colors that vary between sessions. The highlighted colors are selected based on predominant color detection using K-means clustering (Statsmodels 2024), followed by Hue filtering (Chu, Tsuji, and Kato 2014) and manual correction for accuracy. Third, we added bounding boxes to the segmented objects. Fourth, we applied a natural perturbation to the part to be assembled by randomly shifting it within the background box, ensuring it moves by at least 5%. We treated each state from the manual as a positive sample and generated three variant states as negative samples. Then, we constructed the state detection dataset, $D_{T3}=\{(Q_{t},V_{t},A_{t},\tilde{V}_{t},\tilde{A}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}},\tilde{A}_{t}^{{}^{\prime\prime}})\}|_{t=0}^{|D_{T3}|}$, where the query $Q_{t}$ is constructed by filling the query template with the task token “[state]”, the current instruction $I_{t}$, and the previous instruction $I_{t-1}$ as context. The variant states $\{\tilde{V}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}}\}$ are natural pertubations of $V_{t}$. The responses $A_{t}$ and $\{\tilde{A}_{t},\tilde{A}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime\prime}}\}$ indicate the correct and incorrect statuses of the assembly state, respectively.
Dataset Statistics
# Manual # Session # Scene # Object # State
65 397 5,612 4,784 2,716
# Step Overall Identification Assembly
10,428 4,814 5,614
# Sample Overall T1 T2 T3
35,612 5,612 19,136 10,864
Table 1: Statistics of LEGO-VLM dataset.
Experimental Setup
Benchmarks
We benchmark the following nine prevailing VLMs: (1) mPLUG-OWL2 (Ye et al. 2024)replaces attention with a modality adapter in a large language model (LLM) decoder. (2) BLIP2 (Li et al. 2023c)uses a two-stage pre-trained Q-Former between an image encoder and a LLM to bridge the vision-language modality gap. (3) LLaVa (Liu et al. 2023)combines a visual CLIP encoder and a language decoder Vicuna for general-purpose visual language understanding. (4) Qwen-VL (Bai et al. 2023)connects a LLM with a visual encoder using position-aware vision-language adapter towards fine-grained visual understanding. (5) InstructBLIP (Dai et al. 2023)explores general-purpose vision-language instruction tuning based on the pretrained BLIP2. (6) MiniGPT-v2 (Chen et al. 2023)links a frozen ViT visual encoder with Llama-2 via a projection layer, applicable for diverse tasks via task-specific multimodal instructions. (7) Otter (Li et al. 2023a)is tuned on the OpenFlamingo, conditioning the language model on images for multi-modal perception and reasoning. (8) MiniGPT-4 (Zhu et al. 2023)integrates a frozen visual ViT&Q-Former encoder and Vicuna via a projection layer, unlocking advanced multimodal capabilities like GPT-4. (9) GPT-4o(OpenAI 2024)is a widely used commercial model accessible via APIs. (10) Qwen-VL-2.5update Qwen-VL with the latest Qwen LLM (QwenTeam 2024). (11) Gemini-2.5-flash(Comanici et al. 2025)is a thinking model, designed to tackle increasingly complex problems. (12) GLM-4.1-thinking(Hong et al. 2025)is designed to explore the upper limits of reasoning.
Evaluation Metrics
@@
Outcomes
Benchmark Results
@@
mPLUG-OWL2 23.16 29.70 25.23 32.75 3.04 8.46 77.94 93.05 14.23 34.88 36.82 15.00 63.41 35.00
BLIP2 27.85 32.65 29.55 40.35 6.40 12.50 77.35 84.48 34.25 40.57 24.71 28.50 100.00 94.87
LlaVA 32.10 35.88 13.63 42.97 0.99 17.73 54.39 72.32 $\emptyset$ 60.98- 25.04 30.00 99.41 50.00
Qwen-VL 34.84 37.00 29.57 39.08 5.19 13.13 78.56 89.15 25.60 30.08 39.77 39.53 97.99 96.82
InstructBLIP 32.85 32.85- 29.20 36.87 4.91 11.52 79.76 98.16 $\emptyset$ 47.20- 0.00 0.02 1.22 0.00
MiniGPT-v2 33.06 37.52 34.72 32.56 9.81 8.28 84.95 85.91 26.98 25.94 36.76 38.64 60.42 80.72
Otter 11.49 / 12.12 / 1.28 / 72.39 / $\emptyset$ / 35.33 / 75.88 /
MiniGPT-4 34.11 / 15.05 / 1.88 / 87.09 / 30.20 / 37.19 / 67.37 /
GPT-4o 25.81 / 18.67 / 2.00 / 66.67 / 21.68 / 40.54 / 43.06 /
Gemini-2.5-flash 2.31 / 28.28 / 6.21 / 93.12 / 13.16 / 28.81 / 8.72 /
Qwen-VL-2.5 19.13 / 15.34 / 2.70 / 75.55 / 10.63 / 33.85 / 36.70 /
GLM-4.1-thinking 35.73 / 19.44 / 2.31 / 70.64 / 4.40 / 39.61 / 38.53 /
Table 2: Benchmarking vision-language models on LEGO-VLM dataset. The upper part represents open-resourced vision-language models, both without (/wo) and with (/w) PEFT using LoRA. The lower part presents results obtained via API calls to the latest vision-language models. The bold font indicates the highest score in each column. Symbols $\uparrow$ and $\downarrow$ denote that higher and lower values are better, respectively. Symbol “-” indicates the model is not applicable for fine-tuning. Symbol $\emptyset$” denotes a meaningless zero as the model fails to generate output as instructed. The superscripts “”, “”, and “-” indicate an increase, decrease, or inapplicability in the evaluation score after fine-tuning, respectively.
@@
Data Quality Assessment
@@
To ensure the quality of the generated data for tasks T2 and T3, we sampled 100 data samples and conducted quality assessments for each task. We added bounding boxes based on the coordinates and asked three annotators to evaluate the generated data quality, focusing on coordinates, entities, and negative samples, based on the following criteria: (1) Entity disambiguitymeasures how clearly a target entity is relevant to the scene in T2, with scores of 0, 1, and 2 indicating low, medium, and high disambiguation, respectively. (2) Boundary precisionmeasures how accurately a bounding box encloses a target object in T2, with scores of 0, 1, and 2 indicating low, medium, and high precision, respectively. (3) State relevancemeasures whether the generated parts in an image are a relevant variation of the original image in T3. (4) State identifiabilitymeasures whether the generated parts in an image are a recognizable variation of the original image in T3.
Figure 4: Human assessment of data quality.
As shown in Figure 4, the average entity disambiguation and boundary precision scores are 1.74 and 1.25 out of 2 with the Kappa values (Scikit-learn 2024) of 0.60 and 0.61, respectively.
@@
Query: [grounding] After I find 1 transparent mailbox front 2x2. Tell me what to do for the current situation. [Uncaptioned image]
Reference of Assembly Instruction:
@@
Insert the two tabs at the bottom of the mailbox casing into the two small notches on the mailbox front.
mPLUG-OWL2: Place the mailbox front 2x2 next to the block on the table   .
@@
BLIP2: Put the mailbox front 2x2 horizontally on the table, clasp to the back. Put the mailbox back 2x2 vertically on top, slope to the left.
@@
LlaVA: In the image, there is a small Lego-like object with a red liquid inside of it. The object is placed in a blue sky, and it appears to be a 3D model. To create a transparent mailbox, you need to find a transparent mailbox front that is 2x2 in size. Once you find it, you can use it as the base for your transparent mailbox. You can then place the small Lego-like object inside the mailbox to complete the scene.
@@
Qwen-VL: Put the transparent mailbox front 2x2 on the right side of the white arch.
InstructBLIP: Place it horizontally on the table, clasp to the back.
@@
MiniGPT-v2: Here’s what you need to do with the pieces you have collected: 1. Place the transparent mailbox front 2x2 on top of the plate 1x2, aligned with the short side of the plate. Now you have completed the current situation. Let me know if you need help with anything else!
@@
Otter: In view of the reasoning ‘A mailbox is small and square with a see through window.’, the optimal answer should be: build structure.
@@
MiniGPT-4: Attach the transparent mailbox front to the base structure, aligning the tabs with the notches.